OpenStudy (anonymous):

If (x , 3) is on the graph of 2x - 5y =11, what is x?

OpenStudy (anonymous):

just substitute the value 3, into the equation and u can find the valu of x..

OpenStudy (anonymous):

we go to the original equation and rearrange it to make x the subject

OpenStudy (anonymous):

11+5y=2x 11/2+(5/2)y=x

OpenStudy (mathmate):

(x,3) means x is not known, and y=3. As J-D put it, substitute y=3 in the equation of the line and solve for x.

OpenStudy (anonymous):

and we substitute the y into the equation to find x: 5.5+2.5X3=x x=13

OpenStudy (anonymous):

2x-5(3)=11 2x-15=11 2x=15+11 2x=26 \[\frac{ 2x }{ 2? }=\frac{ 26 }{ 2? }\] x= 13 Can someone please check my answer is correct?

OpenStudy (anonymous):

yes it is.....

OpenStudy (anonymous):

yah.thank you. :)

OpenStudy (anonymous):

ur welcome...